a manufacture has 15 years, 8.5 percent mortgage for 100,000 birr on a building. payments are made monthly.

a. find the monthly payment

b. how much interest will be paid?


1
Expert's answer
2022-03-07T18:09:49-0500

m=15

r=8.5%

PV=100000 birr

a)

monthly payment is given as:

"=\frac{PV (\frac{r}{n})(1+\frac{r}{m})^{(mn)}}{(1+\frac{r}{m})^{mn}-1}"


"=\frac{100000(\frac{0.085}{12})(1+\frac{0.085}{12})^{(12×15)}}{(1+\frac{0.085}{12})^{12×15}-1}"


"=984.74\space{birr}" is the monthly payment


b) the total amount to be repaid is

"=984.74\u00d7(12\u00d715)"

"=177253.2\space{birr}"

Interest"=177253.2-100000"

"=77253.2\space{birr}" is the interest paid on the mortgage